Can someone help me find the domain for g(x)=6/2-5x
Answer by nyc_function(2741)

g(x)= 6/2-5x
To domain of g(x) will be the answer(s) for x that do not make the denominator zero.

To find the domain, set 2 - 5x = 0 and solve for x.

2 - 5x = 0

-5x = -2

x = -2/-5

x = 2/5

DOMAIN = ALL REAL NUMBERS except that x CANNOT be 2/5.


Done!

NOTE: If x = 2/5 here, it will create a zero in the denominator, which is undefined since there is no such thing as division by zero.